OpenStudy (anonymous):

A single slit has a width of 12.7 µm. Monochromatic light from a source passes through the slit and a fourth dark fringe (on one side of the central maximum) appears at an angle of 10.0° as measured from the CM. Calculate the wavelength of light used. What colour is this?

OpenStudy (anonymous):

Condition of minimum is \[slit\ width \times sin\ \theta = order \times wavelength\] Put \( \theta = 10.0^o\) , slit width = 12.7 \(\mu m\) and order = 4 .

OpenStudy (anonymous):

Thanks. Got 551 nm which is green.